OpenStudy (anonymous):

write the standward form of the equation of the line through the pair of points (-1,6) and (-8,8). Write the equation using onlt integer coefficients.

OpenStudy (anonymous):

find the slope by y2 - y1/ x2-x1

OpenStudy (anonymous):

8-6/ -8 - (-1)=-2/7.... so y=mx +b y=2/7x + b plug in any of the two points and get b 6= 2/7(-1) + b not sure if fractions are integers or i just did the numbers wrong